OpenStudy (anonymous):

an Individual starts a business and loses $150k in the first yr, $120k in the second yr and $90k in the third yr. If the improvement continues at the same rate. find the profit or loss after 20 yrs. after how many yrs will his loses be juss balanced by his gains

OpenStudy (anonymous):

After 20 years, the business will have 450k profit

OpenStudy (anonymous):

It will take 10 years to balance the losses.

OpenStudy (campbell_st):

find the equation of the line as the data suggests that it is linear  so the break-even is 6 years the profit after 20 years is oops profit is in 100s \[P = 30000 \times 20 - 180000 \] P = 600000 - 180 000 P = $420 000